# TDD Without Tears: Towards Test Case Generation from Requirements through Deep Reinforcement Learning

### v1

@@ 4.4. Baselines
  • GPT-3.5 (175B parameters [^10] GPT-3.5 model size is not reported publicly. The previous version (GPT-3 model(Brown et al., 2020)) has a model parameter size of 175B.(OpenAI, 2023) is one of the most powerful state-of-the-art LLMs developed by OpenAI, which is a transformer-based LLM pre-trained with the next token prediction task on the internet and the third-party licensed data, then finetuned using Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017). We chose the recommended gpt-3.5-turbo endpoint, which was accessed in June 2023.
4.5. Model Experimentation
@@
where $r$ is the reward point, and $p$ is a penalty point, or 0 if not specified. For example, when considering ’Syntax-only’ as a reward function, a reward point of 2 is given if the generated test cases are syntactically correct, otherwise, a penalty point of -2 is given. We run our experiment on two NVIDIA GeForce RTX 3090 GPUs with 24 GB vRAM, an Intel(R) Core(TM) i9-9980XE CPU @ 3.00GHz with 36 core processors, and 64G RAM. The hyperparameter settings for the policy training and the policy optimization are reported in Table 2 and Table 2, respectively.
In total, we conducted an experiment with 7 variations. During the training phase, we set the model to learn to generate multiple assert statements (without the length limit). During the inference phase, the test case generation for each model uses a beam search of 5 and each model is evaluated using the following measure, called TestCaseScore:
(12) $\mathrm{TestCaseScore}=\frac{(2\times CodeCoverage\times MutationScore)}{(CodeCoverage+MutationScore)}$
TestCaseScore is a harmonic mean of the code coverage and the mutation score. After the comprehensive experiment, we find that the PyTester model performs best when it is trained to generate multiple assert statements with a reward function considering both Syntax and Coverage feedback, which we used as a reference setting for our PyTester model for the remainder of the paper.
Table 1. The hyperparameter settings of PyTester during the policy training step.
Table 2. The hyperparameter settings of PyTester during the policy optimization step.
@@
Table 2. The hyperparameter settings of PyTester during the policy optimization step.
5. Results & Discussion
@@
Table 3. (RQ1) The performance of PyTester when compared to state-of-the-art LLMs.
@@
Model ModelParameters Inference Time SyntaxCorrectness Requirement Alignment CodeCoverage Mutation Score
PyTester
770M
¡1 hr
99.42% 84.41% 80.98% 61.45%
Finetuned CodeT5-large
770M
¡1 hr
65.08%
35.56%
34.44%
26.94%
InCoder
6B
250 hr
4.63%
0.03%
0.03%
0.03%
StarCoder
15.5B
40 hr
96.78%
71.56%
69.07%
54.32%
GPT3.5
175B
12 hr
98.90%
40.20%
37.43%
29.00%
@@
Results. Of the test cases generated by PyTester, 99% are syntactically correct, 84% are aligned with the requirements while achieving a code coverage of 80% and a mutation score of 61%, which outperforms all of the studied LLMs. Table 3 presents the performance of PyTester when compared to the state-of-the-art LLMs according to the four evaluation metrics. When comparing PyTester with GPT3.5, our PyTester model is able to generate test cases that are 1% more syntactically correct, 44% more aligned with the requirement, 44% more complete, and 32% more effective in the shortest amount of inference time due to the smallest size of the model. This finding indicates that a small language model like PyTester which is based on CodeT5 does not necessarily perform worse than large language models if they are carefully designed. Instead, by incorporating the domain knowledge of software testing into the architecture of PyTester (i.e., the consideration of test case characteristics in the reward function that is largely ignored by existing LLM models), PyTester is able to perform better than large language models, with a smaller size of the model parameters and a shorter amount of inference time. This highlights the significant novelty of our approach contributing to the research area of Text-to-Testcase generation.
@@
Our proposed DeepRL framework can improve the requirement alignment percentage by 49%, the code coverage by 47%, and the mutation score by 32% when compared with the CodeT5 base model. Table 3 shows that, without the deep reinforcement learning, the CodeT5 base model performs worse than GPT3.5, which is aligned with the common intuition and prior studies (OpenAI, 2023; Guo et al., 2023). On the other hand, when adding deep reinforcement learning to the CodeT5 base model, our experiment confirms that the PyTester model performs better than the GPT3.5 for all evaluation metrics. This finding highlights the importance of our proposed deep reinforcement learning framework for the Text-to-Testcase generation task.
@@
Figure 4 presents an illustrative example of the requirement input and the test cases by PyTester and the baselines. This example depicts a programming task (Problem ID #569) in the APPS dataset. We observe that PyTester can generate test cases that are syntactically correct and align with the requirement while achieving high code coverage and high mutation score. On the other hand,
@@
Figure 4. An example of test cases generated by PyTester and other baseline approaches (depicted from Problem ID #569 of the testing dataset).
@@
Table 4. (RQ2) The performance of our PyTester for the 7 variations of the reward functions. S: Syntax, E: Executability, C: Coverage.
RewardFunction SyntaxCorrectness Requirement Alignment CodeCoverage Mutation Score TestCaseScore
S-only
99.39%
82.72%
79.47%
60.96%
69.00%
E-only
96.26%
75.73%
72.95%
56.34%
63.58%
C-only
98.59%
86.87%
81.96%
56.47%
66.87%
S+E
98.80%
83.49%
80.34%
61.25%
69.51%
S+C
99.42%
84.41%
80.98%
61.45%
69.87%
E+C
94.57%
67.78%
65.37%
51.01%
57.31%
S+E+C
98.80%
81.68%
78.34%
58.76%
67.15%
Results. The top-3 reward functions for PyTester are the Syntax+Coverage, Syntax+Executability, and Syntax-only types of feedback. Table 4 presents the performance of our PyTester for the seven variations of the reward functions. We find that the Syntax+Coverage, Syntax+Executability, and Syntax-only types of the reward function achieve a comparative TestCaseScore of 69.87%, 69.51%, and 69.00%, respectively. On the other hand, without considering the syntax correctness feedback (for the remaining reward function types), the TestCaseScore of PyTester is decreased by 3-13%. This finding highlights the importance of syntax correctness in the reward function—i.e., such feedback will greatly help the agent to generate test cases that are syntactically correct during the policy optimization step. This finding suggests that syntax correctness must be the minimum consideration when designing a reward function for deep reinforcement learning in the text-to-test case generation task.
In addition, we also find that both Syntax+Coverage and Syntax+Executability mostly perform best (i.e., highly correlated). For Syntax+Coverage, the model aims to directly generate syntactically correct test cases that achieve high code coverage, enabling the generated test cases to be executable against the ground truth code. On the other hand, for Syntax+Executability, the model aims to directly generate syntactically correct test cases that are executable against the ground truth code, enabling the generated test cases to achieve high code coverage. Therefore, considering either Syntax+Coverage or Syntax+Executability in the reward function is the best. Nevertheless, the lowest-performing variation of our PyTester still outperforms GPT3.5 and other LLMs.
@@
Figure 5. (RQ3) The common types of runtime errors that occur in the PyTester-generated test cases.
Results. Among the incorrectly generated test cases, PyTester mostly encounters an Assertion Error issue for 10.75%. Despite PyTester performing the best when compared to the studied LLMs, we want to gain deeper insights into the incorrectly generated test cases. As mentioned in Table 3, we find that 84.41% of the test cases generated from PyTester are aligned with the requirements (the blue color in Figure 5), meaning that 15.59% of the generated test cases are not executable, thus incorrect. Thus, we conduct an analysis to investigate what are the common types of runtime errors that occur in the PyTester-generated test cases. As shown in Figure 5, PyTester encounters the following top five common types of errors as follows:
  • Assertion Error (10.75%) is an exception that is raised when an assert statement fails. The AssertionError is raised when the actual output (i.e., the output of the generated assertion) does not match the expected output (i.e., the output from the ground truth code). This finding highlights that as low as 10.75% of the generated test cases are not executable (i.e., are not aligned with the requirement). Therefore, future researchers should consider developing a novel mechanism integrated into the RL framework to improve the requirement alignment.
@@
6. Related Works

### v2

@@ 4.5 Baselines
  • GPT-3.5 (>175B parameters [^10] GPT-3.5 model size is not reported publicly. The previous version (GPT-3 model(Brown et al., 2020)) has a model parameter size of 175B.(OpenAI, 2023) is one of the most powerful state-of-the-art LLMs developed by OpenAI, which is a transformer-based LLM pre-trained with the next token prediction task on the internet and the third-party licensed data, then finetuned using Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017). We chose the recommended gpt-3.5-turbo endpoint, which was accessed in June 2023.
  • Copilot (>12B parameters [^11] Copilot model size is not reported publicly. The previous version (Codex model(Chen et al., 2021)) has a model parameter size of 12B.(GitHub, 2021) is another one of the most powerful state-of-the-art LLMs from GitHub. It was powered by the OpenAI Codex Model and introduced as an “AI pair programmer” in Visual Studio Code. We include Copilot in this study to reflect the performance of a tool in a realistic coding scenario. Access to Copilot was obtained through the following repository. [^12] https://github.com/B00TK1D/copilot-api/tree/main margin:
  • Pynguin (Lukasczyk and Fraser, 2022) is a search-based test case generation approach that produces test cases given a source code. We include Pynguin in this study to reflect the performance of a traditional Code-to-Testcase tool. Notably, due to a technical issue with IO functions, Pynguin is incompatible with the APPS dataset. Nonetheless, we include Pynguin in the discussion section (see Section 6), where we assess our PyTester in the zero-shot setting on the MBPP and HumanEval datasets.
4.6 Model Experimentation
@@
where $r$ is the reward point, and $p$ is a penalty point, or 0 if not specified. For example, when considering ’Syntax-only’ as a reward function, a reward point of 2 is given if the generated test cases are syntactically correct, otherwise, a penalty point of -2 is given.
We run our experiment on two NVIDIA GeForce RTX 3090 GPUs with 24 GB vRAM, an Intel(R) Core(TM) i9-9980XE CPU @ 3.00GHz with 36 core processors, and 64G RAM. The hyperparameter settings for the policy training and the policy optimization are reported in Table 3 and Table 3, respectively.
In total, we experimented with 7 variations. During the training phase, we set the model to learn to generate multiple assert statements (without the length limit). During the inference phase, the test case generation for each model uses a beam search of 5 and each model is evaluated using the following measure, called TestQuality:
$\mathrm{TestQuality}=\frac{(2\times CodeCoverage\times MutationScore)}{(CodeCoverage+MutationScore)}$ (12)
Ideally, high-quality test cases must execute all lines of code in the function (i.e., measured via code coverage), while being able to detect bugs if the code is slightly changed (i.e., measured via mutation score). We define the Test Quality score as a harmonic mean of the code coverage and the mutation score. After the comprehensive experiment, we find that the PyTester model performs best when it is trained to generate multiple assert statements with a reward function considering both Syntax and Coverage feedback, which we used as a reference setting for our PyTester model for the remainder of the paper.
Table 2: The hyperparameter settings of PyTester during the policy training step.
@@
Table 3: The hyperparameter settings of PyTester during the policy optimization step.
@@
5 Experimental Results
@@
Table 4: (RQ1) The performance of PyTester when compared to state-of-the-art LLMs. LLMs without beam search are run 5 times and reported as Mean (SD).
@@
Evaluation Setting Model ModelParameters Inference Time SyntaxCorrectness Passing Test Case CodeCoverage Mutation Score
Raw PyTester 770M $<$1 hr 99.42% 84.41% 80.98% 61.45%
Finetuned CodeT5-large 770M $<$1 hr 65.08% 35.56% 34.44% 26.94%
InCoder 6B 250 hr 4.63% 0.03% 0.03% 0.03%
StarCoder 15.5B 40 hr 96.78% 71.56% 69.07% 54.32%
GPT3.5 >175B 12 hr 93.59% (0.79%) 79.21% (0.67%) 75.10% (0.68%) 54.32% (0.58%)
Copilot >12B 6 hr 59.40% (0.87%) 48.73% (0.65%) 47.11% (0.60%) 33.82% (0.62%)
3-test PyTester 770M $<$1 hr 99.42% 84.44% 80.92% 60.36%
Finetuned CodeT5-large 770M $<$1 hr 96.01% 53.57% 51.58% 40.23%
InCoder 6B 250 hr 19.24% 0.12% 0.11% 0.07%
StarCoder 15.5B 40 hr 99.29% 72.63% 70.02% 52.59%
GPT3.5 >175B 12 hr 93.73% (0.79%) 80.27% (0.78%) 76.07% (0.77%) 54.53% (0.60%)
Copilot >12B 6 hr 94.42% (0.33%) 72.47% (0.22%) 70.00% (0.25%) 53.18% (0.29%)
5-test PyTester 770M $<$1 hr 99.42% 84.41% 80.91% 60.41%
Finetuned CodeT5-large 770M $<$1 hr 91.53% 38.75% 37.50% 29.42%
InCoder 6B 250 hr 12.83% 0.03% 0.03% 0.03%
StarCoder 15.5B 40 hr 98.96% 71.86% 69.34% 54.38%
GPT3.5 >175B 12 hr 93.72% (0.79%) 79.36% (0.69%) 75.23% (0.68%) 54.07% (0.58%)
Copilot >12B 6 hr 90.72% (0.53%) 61.15% (0.54%) 59.12% (0.55%) 41.08% (0.66%)
Filtered by Syntax Correctness PyTester 770M $<$1 hr 99.63% 84.50% 81.06% 60.54%
Finetuned CodeT5-large 770M $<$1 hr 98.93% 38.39% 36.90% 24.86%
InCoder 6B 250 hr 1.90% 0.12% 0.11% 0.05%
StarCoder 15.5B 40 hr 99.97% 71.95% 69.25% 47.13%
GPT3.5 >175B 12 hr 94.55% (0.70%) 79.79% (0.72%) 75.68% (0.71%) 54.76% (0.59%)
Copilot >12B 6 hr 99.52% (0.03%) 59.59% (0.95%) 57.58% (0.93%) 39.56% (0.83%)
@@
Results. Of the test cases generated by PyTester, 99% are syntactically correct, 84% are passed while achieving a code coverage of 80% and a mutation score of 61%, which outperforms all of the studied LLMs. Table 4 presents the performance of PyTester when compared to the state-of-the-art LLMs according to the four evaluation metrics. When comparing PyTester with GPT3.5, our PyTester model is able to generate test cases that are 6% more syntactically correct, 5% more passing, 6% more complete, and 7% more effective in the shortest amount of inference time due to the smallest size of the model. This finding indicates that a small language model like PyTester which is based on CodeT5 does not necessarily perform worse than large language models if they are carefully designed. Instead, by incorporating the domain knowledge of software testing into the architecture of PyTester (i.e., the consideration of test case characteristics in the reward function that is largely ignored by existing LLM models), PyTester is able to perform better than large language models, with a smaller size of the model parameters and a shorter amount of inference time. This highlights the significant novelty of our approach contributing to the research area of Text-to-Testcase generation.
@@
Our proposed DeepRL framework can improve the passing test case percentage by 49%, the code coverage by 47%, and the mutation score by 32% when compared with the CodeT5 base model. Table 4 shows that, without the deep reinforcement learning, the CodeT5 base model performs worse than GPT3.5, which is aligned with the common intuition and prior studies (OpenAI, 2023; Guo et al., 2023). On the other hand, when adding deep reinforcement learning to the CodeT5 base model, our experiment confirms that the PyTester model performs better than the GPT3.5 for all evaluation metrics. This finding highlights the importance of our proposed deep reinforcement learning framework for Text-to-Testcase generation.
@@
Figure 4 presents an illustrative example of the description input and the test cases by PyTester and the baselines. This example depicts a programming task (Problem ID #569) in the APPS dataset. We observe that PyTester can generate test cases that are syntactically correct and align with the description while achieving high code coverage and high mutation score. On the other hand,
@@
Figure 4: An example of test cases generated by PyTester and other baseline approaches (depicted from Problem ID #569 of the testing dataset).
@@
Table 5: (RQ2) The performance of our PyTester for the 7 variations of the reward functions. S: Syntax, E: Executability, C: Coverage.
RewardFunction SyntaxCorrectness Passing Test Case CodeCoverage Mutation Score TestQuality
S-only 99.39% 82.72% 79.47% 60.96% 69.00%
E-only 96.26% 75.73% 72.95% 56.34% 63.58%
C-only 98.59% 86.87% 81.96% 56.47% 66.87%
S+E 98.80% 83.49% 80.34% 61.25% 69.51%
S+C 99.42% 84.41% 80.98% 61.45% 69.87%
E+C 94.57% 67.78% 65.37% 51.01% 57.31%
S+E+C 98.80% 81.68% 78.34% 58.76% 67.15%
Results. The top-3 reward functions for PyTester are the Syntax+Coverage, Syntax+Executability, and Syntax-only types of feedback. Table 5 presents the performance of our PyTester for the seven variations of the reward functions. We find that the Syntax+Coverage, Syntax+Executability, and Syntax-only types of the reward function achieve a comparative TestQualityScore of 69.87%, 69.51%, and 69.00%, respectively. On the other hand, without considering the syntax correctness feedback (for the remaining reward function types), the TestQualityScore of PyTester is decreased by 3-13%. This finding highlights the importance of syntax correctness in the reward function—i.e., such feedback will greatly help the agent to generate test cases that are syntactically correct during the policy optimization step. This finding suggests that syntax correctness must be the minimum consideration when designing a reward function for deep reinforcement learning in the text-to-test case generation task.
In addition, we also find that both Syntax+Coverage and Syntax+Executability mostly perform best (i.e., highly correlated). For Syntax+Coverage, the model aims to directly generate syntactically correct test cases that achieve high code coverage, enabling the generated test cases to be executable against the ground truth code. On the other hand, for Syntax+Executability, the model aims to directly generate syntactically correct test cases that are executable against the ground truth code, enabling the generated test cases to achieve high code coverage. Therefore, considering either Syntax+Coverage or Syntax+Executability in the reward function is the best. Nevertheless, the lowest-performing variation of our PyTester still outperforms most of the LLMs.
@@
Figure 5: (RQ3) The common types of runtime errors that occur in the PyTester-generated test cases.
Results. Among the incorrectly generated test cases, PyTester mostly encounters an Assertion Error issue for 10.75%. Despite PyTester performing the best when compared to the studied LLMs, we want to gain deeper insights into the incorrectly generated test cases. As mentioned in Table 4, we find that 84.41% of the test cases generated from PyTester are passed (the blue color in Figure 5), meaning that 15.59% of the generated test cases are not executable, thus incorrect. Thus, we conduct an analysis to investigate what are the common types of runtime errors that occur in the PyTester-generated test cases. As shown in Figure 5, PyTester encounters the following top five common types of errors as follows:
  • Assertion Error (10.75%) is an exception that is raised when an assert statement fails. The AssertionError is raised when the actual output (i.e., the output of the generated assertion) does not match the expected output (i.e., the output from the ground truth code). This finding highlights that as low as 10.75% of the generated test cases are not passed the ground truth code. Therefore, future researchers should consider developing a novel mechanism integrated into the RL framework to improve the passing test case.
@@
margin:
6 Discussion
In this section, we discuss the additional evaluation of PyTester in the zero-shot setting on the HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) datasets. Notably, the HumanEval and MBPP datasets are relatively small, with sample sizes of $n$=164 and $n$=257, respectively.
On the HumanEval dataset, the raw output of PyTester ranks the second place in generating executable test cases with high code coverage, following only GPT-3.5. Table 7 shows the results of PyTester compared to state-of-the-art LLMs and Pynguin on the HumanEval dataset. Although performing in the zero-shot setting, the results show that PyTester can generate syntactically correct, executable test cases with high code coverage comparable to LLMs such as StarCoder and Copilot. In fact, our PyTester model achieves the second rank in the raw evaluation, following only the GPT-3.5 model. It is important to note that albeit providing the necessary information (i.e., description text, code, and test cases), the HumanEval dataset typically provides a concise docstring for a function, different to the APPS dataset (i.e., the training dataset) that contains a lengthy, detailed description for a function. Therefore, this difference presents limitations that can impact the performance of PyTester.
On the MBPP dataset, PyTester performance is comparable to the StarCoder Model. Table 7 shows the results of PyTester compared to state-of-the-art LLMs and Pynguin on the MBPP dataset. Similar to the HumanEval dataset, the MBPP dataset has a unique nature of the text description, which is significantly different to the APPS dataset (i.e., the training dataset). Specifically, the MBPP dataset provides only a single-line description of a function, different to the APPS dataset that contains a lengthy, detailed description for a function. Therefore, this characteristic significantly impacts the performance of our PyTester that trained on a detailed description. However, the results show that our PyTester can still perform better than the Finetune CodeT5-large and the InCoder models, and comparable if not better than the significantly larger model such as the StarCoder model.
Table 6: The performance of PyTester in the zero-shot setting on HumanEval when compared to state-of-the-art LLMs and Pynguin.
Evaluation Setting Model ModelParameters SyntaxCorrectness Passing Test Case CodeCoverage MutationScore
Raw PyTester 770M 99.39% 83.54% 80.09% 48.74%
Finetuned CodeT5-large 770M 96.34% 68.90% 68.02% 49.80%
InCoder 6B 12.20% 4.27% 4.16% 4.10%
StarCoder 15.5B 85.37% 78.05% 77.25% 63.75%
GPT3.5 >175B 99.39% (0.61%) 89.51% (0.80%) 88.36% (0.96%) 69.59% (0.76%)
Copilot >12B 83.78% (0.92%) 79.51% (0.70%) 79.00% (0.69%) 63.50% (0.49%)
Pynguin - 86.59% 83.54% 82.87% 25.59%
3-test PyTester 770M 99.39% 83.54% 80.11% 48.74%
Finetuned CodeT5-large 770M 100.00% 73.17% 71.67% 50.58%
InCoder 6B 54.88% 14.63% 14.28% 11.80%
StarCoder 15.5B 100.00% 85.37% 83.99% 65.82%
GPT3.5 >175B 99.51% (0.51%) 92.80% (1.00%) 91.19% (1.27%) 70.62% (1.15%)
Copilot >12B 98.78% (0.00%) 89.63% (0.43%) 88.62% (0.43%) 70.52% (0.25%)
Pynguin - 87.80% 85.98% 85.21% 23.61%
5-test PyTester 770M 99.39% 83.54% 80.11% 48.74%
Finetuned CodeT5-large 770M 99.39% 69.51% 68.63% 50.28%
InCoder 6B 46.34% 7.93% 7.80% 7.17%
StarCoder 15.5B 100.00% 80.49% 79.65% 64.35%
GPT3.5 >175B 99.51% (0.51%) 89.87% (1.26%) 88.68% (1.42%) 69.62% (0.89%)
Copilot >12B 97.32% (0.33%) 85.12% (0.33%) 84.51% (0.33%) 68.07% (0.23%)
Pynguin - 85.37% 84.15% 83.43% 21.45%
Filtered by Syntax Correctness PyTester 770M 98.78% 83.54% 80.11% 48.74%
Finetuned CodeT5-large 770M 100.00% 68.90% 68.04% 49.80%
InCoder 6B 35.37% 4.27% 4.16% 4.10%
StarCoder 15.5B 100.00% 82.32% 81.54% 67.30%
GPT3.5 >175B 99.76% (0.33%) 89.75% (0.80%) 88.60% (0.98%) 69.81% (0.95%)
Copilot >12B 99.39% (0.00%) 83.54% (0.61%) 83.01% (0.61%) 66.50% (0.52%)
Pynguin - 86.59% 83.54% 82.87% 25.59%
Table 7: The performance of PyTester in the zero-shot setting on MBPP when compared to state-of-the-art LLMs and Pynguin.
Evaluation Setting Model ModelParameters SyntaxCorrectness Passing Test Case CodeCoverage Mutation Score
Raw PyTester 770M 90.66% 15.18% 14.95% 5.74%
Finetuned CodeT5-large 770M 88.72% 2.72% 2.66% 0.70%
InCoder 6B 27.63% 2.33% 2.31% 1.48%
StarCoder 15.5B 43.19% 9.73% 9.62% 5.20%
GPT3.5 >175B 98.05% (1.48%) 37.04% (1.54%) 36.78% (1.23%) 21.97% (1.30%)
Copilot >12B 61.56% (1.39%) 28.56% (1.05%) 28.54% (1.05%) 16.46% (0.44%)
Pynguin - 89.49% 79.77% 75.86% 20.71%
3-test PyTester 770M 90.66% 15.18% 14.95% 5.74%
Finetuned CodeT5-large 770M 98.44% 4.67% 4.61% 1.87%
InCoder 6B 61.87% 2.72% 2.69% 1.48%
StarCoder 15.5B 87.94% 17.51% 17.42% 9.41%
GPT3.5 >175B 98.44% (1.46%) 38.60% (1.68%) 38.24% (1.50%) 22.76% (0.89%)
Copilot >12B 98.52% (0.17%) 38.68% (0.59%) 38.58% (0.59%) 21.61% (0.43%)
Pynguin - 91.44% 80.16% 76.67% 19.02%
5-test PyTester 770M 90.66% 15.18% 14.95% 5.74%
Finetuned CodeT5-large 770M 95.33% 3.11% 3.05% 1.09%
InCoder 6B 61.87% 2.33% 2.31% 1.48%
StarCoder 15.5B 78.99% 12.84% 12.73% 6.57%
GPT3.5 >175B 98.05% (1.48%) 37.04% (1.54%) 36.78% (1.23%) 21.97% (1.20%)
Copilot >12B 95.85% (0.39%) 34.71% (0.70%) 34.67% (0.70%) 20.58% (0.42%)
Pynguin - 90.66% 80.16% 77.00% 21.71%
Filtered by Syntax Correctness PyTester 770M 89.49% 15.18% 14.95% 5.74%
Finetuned CodeT5-large 770M 99.22% 2.72% 2.66% 0.70%
InCoder 6B 32.30% 2.33% 2.31% 1.48%
StarCoder 15.5B 91.44% 15.56% 15.46% 7.74%
GPT3.5 >175B 98.36% (0.51%) 37.04% (1.54%) 36.78% (1.23%) 21.90% (1.09%)
Copilot >12B 100.00% (0.00%) 35.25% (0.65%) 35.23% (0.65%) 19.80% (0.28%)
Pynguin - 89.49% 79.77% 75.86% 20.71%
7 Related Works